OpenStudy (sambam):

_____ 4√(x^2) Simplify and leave in radical form, Please help

ganeshie8 (ganeshie8):

\(\huge 4\sqrt{x^2}\)

ganeshie8 (ganeshie8):

square root of x^2 = ?

OpenStudy (sambam):

What is that gonna be?

ganeshie8 (ganeshie8):

square root of x^2 = x

ganeshie8 (ganeshie8):

\(\huge 4\sqrt{x^2}\) \(\huge 4x\)

OpenStudy (sambam):

Could you help me with this one? ________ 8√x^2 Y^2

ganeshie8 (ganeshie8):

sure, let me ask u a q first

ganeshie8 (ganeshie8):

that 8 is outside radical, or inside ?

ganeshie8 (ganeshie8):